Simplify trigonometric expression as a single real number
1.(tan Ө)(1/sec Ө)

2.csc(-Ө)/sec(-Ө)

3. (-sin(-Ө))(sec π/2 -Ө))



Answer by stanbon(75887)

Simplify trigonometric expression as a single real number
1.(tan Ө)(1/sec Ө) = (sin/cos)(cos) = sin(theta)
--------------------------------
2.csc(-Ө)/sec(-Ө) = [-csc(theta)]/[sec(theta)]= -cos/sin = -cot(theta)
--------------------------------
3. (-sin(-Ө))(sec (π/2 -Ө)) = (sin(theta)*(csc(theta)) = 1
=====================
Cheers,
Stan H.
====================
--